OpenStudy (queen02):

Find the maximum and minimum values by evaluating the equation, C=4x-3 at the critical points if the restrictions include the following: x≥0 y≥4 x+y≤15 Answers: a. Minimum: -32; Maximum: 45 b. Min: -45; Max:32 c. Min: -45; Max: -32 d. min: 32; Max: 45 @MrNood Can You help?

OpenStudy (queen02):

Please Help! Pleaseee

OpenStudy (queen02):

@MrNood can you help?

OpenStudy (anonymous):



OpenStudy (anonymous):

find C at A,B,C

OpenStudy (anonymous):

in problem C=4x-3y

OpenStudy (anonymous):

at(0,4) C=4*0-3*4=-1 at (0,15) C=4*0-3*15=-45 at (11,4) C=4*11-3*4=44-12=32 can you see which is maximum and which is minimum?